OpenStudy (anonymous):

solve the initial value problem using laplace transform w''(t)-2w'(t)+w=6t-2 when w(-1)=3, w'(-1)=7 I know you have to use y(t)=w(t-1), but i don't know if i have is right! my solution that i found was w(t)= 10+2t-4e^(t+1)+5te^(t+1)

OpenStudy (anonymous):

\[w''(t)-2w'(t)+w=6t-2\] Let \(x=t+1\), so \(t=x-1\). \[w''(x-1)-2w'(x-1)+w=6(x-1)-2\] Let \(w(t)=w(x-1)=y(x)\), so that \(w'(t)=y'(x)\) and \(w''(t)=y''(x)\). \[y''(x)-2y'(x)+y=6x-8\] with initial conditions \[w(-1)=y(-1+1)=y(0)=3\\ w'(-1)=y'(-1+1)=y'(0)=7\] Taking the LT, you have \[\begin{align*} \mathcal{L}\{y''(x)-2y'(x)+y\}&=\mathcal{L}\{6x-8\}\\ (s^2\mathcal{L}\{y\}-sy(0)-y'(0))-2(s\mathcal{L}\{y\}-y(0))+\mathcal{L}\{y\}&=\frac{6}{s^2}-\frac{8}{s}\\ (s^2-2s+1)\mathcal{L}\{y\}-3s-1&=\frac{6}{s^2}-\frac{8}{s}\\ \mathcal{L}\{y\}&=\frac{6-8s+s^2+3s^3}{s^2(s^2-2s+1)} \end{align*}\] Is this what you did?

OpenStudy (loser66):

@SithsAndGiggles I don't understand why we have to convert to y ? Why don't we solve directly from the original problem?

OpenStudy (anonymous):

See example 4 in the link: http://tutorial.math.lamar.edu/Classes/DE/IVPWithLaplace.aspx

OpenStudy (loser66):

I got it, thank you very much :)

OpenStudy (anonymous):

Sorry, slight mistake: \(x-1=t\) Then \(w(t)=w(x-1)=y(x)=y(t+1)\), which gives \(w(-1)=y(0)\).

OpenStudy (anonymous):

yes i have that work! but, when i break up the denominator into partial fractions to simplify, i must be doing something wrong..

OpenStudy (anonymous):

when i break into partial fractions i get: =\[\frac{ 4 }{ s} +\frac{ 6 }{ s ^{2} }-\frac{ 1 }{ s-1 }+\frac{ 2}{ (s-1)^2 }\]

OpenStudy (anonymous):

is that right? my final answer is currently, w(t)= 12+6t-e^(t+1) +2te^(t+1)

OpenStudy (anonymous):

Hmm, that doesn't look completely correct... Here's the partial fraction decomposition: \[\frac{3s^3+s^2-8s+6}{s^2(s-1)^2}=\frac{a}{s}+\frac{b}{s^2}+\frac{c}{s-1}+\frac{d}{(s-1)^2}\] which gives \[3s^3+s^2-8s+6=as(s-1)^2+b(s-1)^2+cs^2(s-1)+ds^2\] or \[3s^3+s^2-8s+6=(a+c)s^3+(-2a+b-c+d)s^2+(a-2b)s+b\] So you have \[\begin{cases}a+c=3\\-2a+b-c+d=1\\a-2b=-8\\b=6\end{cases}~~\Rightarrow~~a=4,~b=6,~c=-1,~d=2\] So your decomposition is right - I just wanted to check for myself. Now, you want to get the following: \[y(x)=\mathcal{L}^{-1}\left\{\frac{4}{s}+\frac{6}{s^2}-\frac{1}{s-1}+\frac{2}{(s-1)^2}\right\}\] Taking the inverse transforms, you have \[y(x)=4+6x-e^x+2xe^x\] Back-substituting with \(x=t+1\), you ahve \[\begin{align*}y(t+1)&=w(t)\\ &=4+6(t+1)-e^{t+1}+2(t+1)e^{t+1}\\ &=4+6t+6-e^{t+1}+2te^{t+1}+2e^{t+1}\\ &=10+6t+e^{t+1}+2te^{t+1}\end{align*}\] You weren't too far off.

OpenStudy (anonymous):

thank you very much! I understand it now, great instructions!

OpenStudy (anonymous):

yw